OpenStudy (marmar10):

Aluminum has a specific heat of 0.215 cal/(g⋅∘C). When 25.6 cal of heat is added to 18.4 g of aluminum at 22.0 ∘C, what is the final temperature of the aluminum?

OpenStudy (whpalmer4):

\[\Delta Q = m c \Delta T\] solve that for \(\Delta T\) and use it to find the change in temperature, which you then add to the starting temperature to find the final temperature.

OpenStudy (marmar10):

34 C?

OpenStudy (marmar10):

it says answer in 3 sig figs and i put 34.0... still wrong hmmm

OpenStudy (whpalmer4):

that's not what I get... show me your work?

OpenStudy (marmar10):

\[25.6= 18.4*.215*Tf-22\] \[47.6= 3.956Tf\] Tf=12.032 12.032+22=34.0

OpenStudy (whpalmer4):

no...get that -22 out of there.

OpenStudy (whpalmer4):

change in heat = mass * c * change in temperature. starting temperature does not matter

OpenStudy (marmar10):

ohhhh

OpenStudy (marmar10):

so i got 28.5 now

OpenStudy (whpalmer4):

still not what I'm getting, but me check here... \[\Delta Q = mc\Delta T\]\[\Delta T = \frac{\Delta Q}{m c} = \frac{25.6 \text{ cal}}{18.4 \text{ g} * 0.215 \text{ cal/g }{^\circ \text{C}}}\]

OpenStudy (whpalmer4):

sorry, that is what I'm getting :-)

OpenStudy (whpalmer4):

was looking at the wrong number on my scratch paper

OpenStudy (marmar10):

awesome thank you!